A cell supplies a current of 0.9A through a 2 ohams resistor. The same cell supplies a current of 0.3 A through a 7 ohams resistor. The internal resistance of the cell is

Grade:12

1 Answers

Soumya Ranjan Mohanty
6 years ago
Dear student,
we know, E=i(R+r)
when, R=2 ohm, i=0.9A so E=0.9(2+r) --(1)
when R=7 ohm,i=0.3 A so E=0.3(7+r)--(2)
equating (1) and (2):
0.9(2+r)=0.3(7+r)
solving =>6+3r = 7+r
we get, r=0.5 ohms (ans.)
Hope it helped.
Thanks,
Soumya Ranjan Mohanty.